### v1

# Gravitational quasinormal modes of Dymnikova black holes

Abstract
We investigate gravitational quasinormal modes of the Dymnikova black hole, a regular spacetime in which the central singularity is replaced by a de Sitter core. This geometry, originally proposed as a phenomenological model, also arises naturally in the framework of Asymptotically Safe gravity, where quantum corrections lead to a scale-dependent modification of the Schwarzschild solution. Focusing on axial gravitational perturbations, we compute the dominant quasinormal frequencies using the WKB method with Padé approximants and verify the results with time-domain integration. We find that the introduction of the quantum parameter $l_{\rm cr}$ leads to systematic deviations from the Schwarzschild spectrum: the real oscillation frequency decreases as $l_{\rm cr}$ increases, while the damping rate also becomes smaller, implying longer-lived modes. In the limit of large $l_{\rm cr}$, the quasinormal spectrum smoothly approaches the Schwarzschild case. These results suggest that even though the corrections are localized near the horizon, they leave imprints in the gravitational-wave ringdown which may become accessible to observation with future high-precision detectors.
I Introduction
One of the most powerful tools of black-hole physics comes from the study of quasinormal modes (QNMs) Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011); Bolokhov and Skvortsova (2025a), which govern the damped oscillations of spacetime following a perturbation. These modes encode key information about the underlying geometry and dominate the ringdown phase of gravitational-wave signals detected by LIGO, Virgo, and KAGRA Abbott et al. (2016a, 2017, 2020, b). The spectrum of gravitational perturbations provides a unique fingerprint of the compact object, enabling tests of general relativity in the strong-field regime and offering the possibility of constraining or ruling out alternative models of gravity and exotic compact objects. For this reason, accurate determination of gravitational QNMs has become a central task in both theoretical and observational astrophysics.
Alongside these observational advances, increasing attention has been directed toward regular black holes—geometries in which curvature singularities are avoided while preserving the main features of standard black-hole spacetimes. Such models provide a phenomenological window into possible effects of quantum gravity and serve as effective descriptions of singularity resolution. They have been studied from multiple perspectives, including energy conditions, stability, and thermodynamics, and have been confronted with astrophysical data in the context of shadows, gravitational lensing, and quasinormal ringing. In particular, QNM spectra of regular black holes often display deviations from the Schwarzschild case that can, in principle, leave observable imprints in gravitational-wave signals.
Quantum effects are likely to leave their imprint not only on the geometry but also on the behavior of test particles, potentially altering aspects of their dynamics Turimov et al. (2025). Moreover, in scenarios where matter fields interact with the background in a non-minimal fashion, or when analyzing associated radiation phenomena, the role of boundary conditions may also acquire particular meaning Derkach et al. (2025); Rakhmanov et al. (2025).
Among the various proposals, the solution introduced by Dymnikova Dymnikova (1992) stands out as an elegant example. It describes a static, spherically symmetric, asymptotically flat geometry in which the central singularity is replaced by a de Sitter core, leading to finite curvature invariants everywhere. While originally formulated as a phenomenological construction, the same functional form has more recently reappeared in the framework of Asymptotic Safety, where it can be derived from renormalization-group improvements tied to curvature invariants Platania (2019). Thus, the Dymnikova spacetime provides not only one of the earliest regular black-hole models, but also a concrete realization of quantum-gravity inspired corrections. This dual interpretation makes it a natural and timely candidate for detailed studies of gravitational perturbations and their observational signatures.
QNMs of test fields in Dymnikova black holes have been investigated in four and higher dimensions in several recent works Konoplya and Zhidenko (2024a); Macêdo et al. (2024); Konoplya et al. (2023). However, to the best of our knowledge, no analogous analysis has yet been carried out for the most important case of gravitational perturbations within the framework of Asymptotically Safe gravity. In this paper we address this gap by computing the dominant QNMs of gravitational perturbations in the Dymnikova background. It is worth noting that grey-body factors and absorption cross-sections for gravitational perturbations of the Dymnikova solution were recently computed in Dubinsky (2025a). However, that analysis did not provide any results for the quasinormal frequencies.
It is worth mentioning that the literature on QNMs of quantum-corrected black holes is vast; therefore, we cite here only a few works Konoplya and Zhidenko (2020); Chen et al. (2023); Baruah et al. (2023); Moreira et al. (2023); Fu et al. (2024); Heidari et al. (2023); Skvortsova (2025); Kokkotas et al. (2017); Abdalla et al. (2005); Skvortsova (2024a); Bonanno et al. (2025); Bolokhov et al. (2025); Bolokhov and Skvortsova (2025b); Malik (2025a, b), referring the reader to the references therein for further studies.
The paper is organized as follows. In Sec. II we introduce the Dymnikova black-hole geometry in the context of Asymptotic Safety. Sec. III is devoted to the analysis of axial gravitational perturbations and the derivation of the effective potential. In Sec. IV we describe the numerical techniques used for calculating QNMs, including the WKB method with Padé approximants and the time-domain integration scheme. The obtained results for the quasinormal spectrum are presented and discussed in Sec. IV. Finally, in Sec. V we summarize our findings and outline possible directions for future research.
II The Dymnikova Black Hole Geometry
The quest for singularity-free black-hole solutions has long motivated the construction of effective geometries that remain regular throughout spacetime. A particularly influential example was proposed by Dymnikova Dymnikova (1992), who suggested replacing the central singularity of the Schwarzschild solution with a de Sitter core while preserving asymptotic flatness. This is achieved through a stress–energy distribution that formally violates some of the classical energy conditions but can be regarded as an effective description of quantum backreaction. The resulting geometry is a static, spherically symmetric, and asymptotically flat spacetime in which all curvature invariants remain finite.
The line element retains the usual Schwarzschild form,
$ds^{2}=-f(r)\,dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}\bigl(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\bigr),$ (1)
but with a modified lapse function
$f(r)=1-\frac{2M(r)}{r},$ (2)
where the mass function interpolates between zero at the origin and the asymptotic ADM mass $M$ at infinity. Dymnikova’s original proposal was
$M(r)=M\Bigl(1-e^{-r^{3}/r_{0}^{3}}\Bigr),$ (3)
with $r_{0}$ characterizing the size of the de Sitter core. At small $r$, the geometry tends to
$f(r)\approx 1-\frac{r^{2}}{r_{0}^{2}},$
describing a regular de Sitter interior, while for large $r$ one recovers the Schwarzschild behavior $f(r)\approx 1-2M/r$. The spacetime thus interpolates smoothly between two familiar limits while remaining free of singularities.
Although originally introduced as a phenomenological model, essentially the same functional form was later derived within the framework of Asymptotic Safety. In this approach, the Schwarzschild lapse function
$f(r)=1-\frac{2M}{r},$ (4)
is modified by allowing Newton’s constant to run with the radial scale. A convenient parametrization is
$G(r)=\frac{G_{0}}{1+g_{\ast}^{-1}G_{0}k^{2}(r)},$ (5)
where $G_{0}$ is the classical Newton coupling, $g_{\ast}$ denotes the fixed-point value, and $k(r)$ is an effective cutoff scale satisfying $k(r)\!\to\!0$ as $r\!\to\!\infty$, thus ensuring the recovery of classical Schwarzschild behavior at large distances. Substituting $G(r)$ into the metric leads to
$f(r)=1-\frac{2M}{r}\,\frac{G(r)}{G_{0}}.$ (6)
The modified lapse can be reinterpreted as arising from an effective energy–momentum tensor,
$T_{\mu\nu}=pg_{\mu\nu}+(\rho+p)(l_{\mu}n_{\nu}+l_{\nu}n_{\mu}),$ (7)
where $l_{\mu}n^{\mu}=-1$. The effective energy density and pressure are generated by the radial dependence of $G(r)$:
$p=-\frac{MG^{\prime\prime}(r)}{8\pi r\,G(r)},\qquad\rho=\frac{MG^{\prime}(r)}{4\pi r^{2}G(r)}.$ (8)
This stress tensor can be interpreted as the imprint of quantum vacuum polarization: deviations of $G(r)$ from constancy induce effective matter sources that regularize the geometry.
A self-consistent solution is obtained through an iterative renormalization-group improvement: starting with the classical Schwarzschild case ($G(r)=G_{0}$), one updates the cutoff $k(r)$ as a functional of the energy density generated at the previous step. In the continuum limit, this iterative process converges to a closed analytic expression for the lapse, which coincides with the Dymnikova form Platania (2019):
$f(r)=1-\frac{2M}{r}\left(1-e^{-r^{3}/(2l_{\rm cr}^{2}M)}\right).$ (9)
Here $l_{\rm cr}$ is a critical length scale marking the onset of quantum corrections. For $l_{\rm cr}\!\to\!0$ the Schwarzschild metric is recovered, while finite $l_{\rm cr}$ introduces a de Sitter-like core. The existence of an event horizon requires
$l_{\rm cr}\lesssim 1.138\,M,$
with larger values of $l_{\rm cr}$ corresponding instead to horizonless, compact, but nonsingular configurations. From now on, we will use units $M=1$. Thus, $\omega$ implies $\omega M$ in dimensionless units.
The Dymnikova solution enjoys a dual origin: historically as a phenomenological model of singularity resolution and, more recently, as a concrete realization of renormalization-group improvements in Asymptotic Safety. This dual perspective underscores its relevance as a minimal yet robust model of a regular black hole. Furthermore, higher-dimensional generalizations of the Dymnikova geometry have been shown to arise naturally in theories with higher-curvature corrections Konoplya and Zhidenko (2024a), further highlighting its role as a testing ground for semiclassical and quantum-gravity inspired modifications of black-hole physics.
III Gravitational perturbations
Figure 1: Effective potential as a function of the tortoise coordinate $r^{*}$ for $\ell=2$, $l_{cr}=0.1$ (blue), $l_{cr}=1$ (red), $l_{cr}=1.137$ (green).
Figure 2: Effective potential as a function of the tortoise coordinate $r^{*}$ for $\ell=3$, $l_{cr}=0.1$ (blue), $l_{cr}=1$ (red), $l_{cr}=1.137$ (green).
Studying gravitational perturbations in spacetimes that incorporate quantum corrections is more subtle than in the classical case. The challenge arises because such geometries, including the Dymnikova black hole, are not derived as exact solutions of Einstein’s equations with a specified matter source, but rather emerge from effective constructions motivated by quantum-gravity considerations. As a result, a straightforward derivation of perturbation equations is not possible. A practical resolution was suggested in the loop-inspired framework of Ashtekar, Olmedo, and Singh Ashtekar et al. (2018a, b), where the background geometry can be reinterpreted within Einstein’s theory if one introduces an effective anisotropic fluid stress tensor that captures the relevant quantum corrections. This reformulation allows the machinery of black-hole perturbation theory to be applied in a consistent manner.
Following this idea, Bouhmadi-López and collaborators Bouhmadi-López et al. (2020); Konoplya and Stashko (2025a) analyzed axial perturbations under the assumption that fluctuations aligned with the anisotropy do not contribute in the axial sector. This simplification is reminiscent of the treatment in Bronnikov et al. (2012), where the decoupling of scalar and axial gravitational modes was exploited in scalar–tensor models (see also Chen and Chen (2019)). Although such assumptions necessarily restrict the analysis, they are widely adopted in related contexts and are generally expected to provide reliable results for small deviations from Schwarzschild geometry Konoplya and Stashko (2025b, a); Dubinsky (2025a); Bolokhov and Skvortsova (2025c).
Under these conditions, the perturbation equations can be reduced to a master wave equation of Schrödinger type,
$\frac{d^{2}\Psi}{dr_{*}^{2}}+\bigl(\omega^{2}-V(r)\bigr)\Psi=0,$ (10)
with the tortoise coordinate defined by $dr^{*}/dr=1/f(r)$. The effective axial potential takes the form
$V(r)=f(r)\left[\frac{(\ell+2)(\ell-1)}{r^{2}}-\frac{f^{\prime}(r)}{r}+\frac{2f(r)}{r^{2}}\right],$ (11)
where $\ell=2,3,4,\ldots$ labels the multipole number. Unlike the standard Regge–Wheeler potential, this expression incorporates contributions from perturbations of the effective fluid stress tensor. Consequently, it deviates from the vacuum Schwarzschild case even when $f(r)$ reduces to the Schwarzschild form.
Figures 1 and 2 show the effective potentials for axial modes in the Dymnikova background. When the quantum parameter $l_{\rm cr}$ is very small (e.g., $l_{\rm cr}=0.01M$), the potential nearly coincides with the Schwarzschild one except in the immediate vicinity of the horizon. The deformation is highly localized, and the potential rapidly merges with the Schwarzschild profile at moderate radii. At the same time, the location of the event horizon shifts inward as $l_{\rm cr}$ grows, leading to a decrease in the horizon radius compared to the Schwarzschild value. These features indicate that the quantum-inspired modifications are concentrated in the near-horizon region, leaving the asymptotic structure essentially unaffected.
Analogous simplifications have been applied in several earlier studies of black-hole perturbations Berti et al. (2003); Kokkotas (1993); Konoplya and Zhidenko (2007). While such approaches neglect certain couplings that might alter the detailed spectrum, they capture the leading behavior whenever the corrections relative to Schwarzschild are small. This perspective fits well with the perturbative character of quantum-gravity modifications, which are expected to act only as moderate deformations of the classical background.
IV Quasinormal modes
For black holes, $V(r)$ has a single peak at $r=r_{0}$ (or $r_{*}=r_{*0}$), vanishes exponentially as $r_{*}\to$-$\infty$ (event horizon), and falls off as $r_{*}\to+\infty$ (spatial infinity). QNMs are defined by the purely ingoing/outgoing asymptotics
$\Psi\propto e^{-i\omega r_{*}},\quad r_{*}\to-\infty\quad(\text{horizon}),$ (12)
$\Psi\propto e^{+i\omega r_{*}},\quad r_{*}\to+\infty\quad(\text{infinity})\,.$ (13)
These conditions select a discrete set of complex frequencies $\omega=\omega_{R}-i\omega_{I}$ with $\omega_{I}>0$. Following the Schutz–Will/Iyer–Will approach Schutz and Will (1985); Iyer and Will (1987) refined to higher orders Konoplya (2003); Matyjasek and Opala (2017); Konoplya et al. (2019), define
$Q(r_{*})\equiv\omega^{2}-V(r_{*}),\quad V_{0}\equiv V(r_{*0}),\quad V_{0}^{(k)}\equiv\left.\frac{d^{k}V}{dr_{*}^{k}}\right|_{r_{*0}}.$ (14)
Expanding around the peak and matching WKB solutions across the turning points yields the $N$-th order quantization condition
$\frac{i\bigl(\omega^{2}-V_{0}\bigr)}{\sqrt{-2V_{0}^{(2))}}}-\sum_{k=2}^{N}\Lambda_{k}\!\left(\{V_{0}^{(j)}\},n\right)=n+\frac{1}{2}\,,n=0,1,2,\ldots,$ (15)
where primes denote $r_{*}$-derivatives, and the $\Lambda_{k}$ are explicit polynomials in $V_{0}^{(j)}$ ($2<j\leq 2k$) and $n$, divided by the corresponding powers of $V_{0}^{(2)}$ (see Iyer and Will (1987); Konoplya (2003); Matyjasek and Opala (2017); Konoplya et al. (2019) for closed forms up to high orders).
Over the past decades, the WKB approach and its refinements have become one of the most widely employed semi-analytical tools for studying black-hole spectra. Its applications extend far beyond the Schwarzschild case, encompassing higher-curvature theories, quantum-corrected geometries, and a variety of exotic compact objects. Numerous examples of such developments can be found in Matyjasek (2021); Pedrotti and Calzà (2025); Konoplya and Zinhailo (2020); Mahdavian Yekta et al. (2021); Zhang et al. (2025); Antonelli et al. (2025); Lütfüoğlu (2025a); Hamil and Lütfüoğlu (2025); Miyachi et al. (2025); Matyjasek and Opala (2017); Konoplya and Abdalla (2005); Konoplya (2021); Konoplya and Zhidenko (2023), which collectively demonstrate both the flexibility and robustness of the method across different gravitational settings.
In the present analysis, our focus is on extracting the two lowest-lying quasinormal frequencies, $\omega_{0}$ and $\omega_{1}$. For this purpose, we employ the sixth or higher-order WKB formalism supplemented by Padé approximants with indices $\tilde{m}=\tilde{n}=3$ Konoplya et al. (2019). This refinement is known to provide optimal accuracy in many situations, as supported by a wide range of recent investigations Dubinsky (2025b); Skvortsova (2024b); Bronnikov and Konoplya (2020); Konoplya (2020); Bolokhov (2024a); Dubinsky (2025c); Skvortsova (2024c); Dubinsky and Zinhailo (2024); Malik (2024a); Churilova et al. (2021); Lütfüoğlu (2025b); Dubinsky (2025d); Malik (2025c); Dubinsky (2025e); Skvortsova (2024d); Bolokhov (2024b); Dubinsky and Zinhailo (2025). The convergence of results across these diverse studies highlights the reliability of the WKB–Padé scheme for accurately capturing the dominant part of the quasinormal spectrum.
It is convenient to introduce a bookkeeping parameter $\epsilon$ (set to 1 at the end) and write
$\frac{i\bigl(\omega^{2}-V_{0}\bigr)}{\epsilon\sqrt{-2V_{0}^{(2)}}}-\sum_{k=2}^{N}\epsilon^{k-1}\Lambda_{k}=n+\frac{1}{2}\,$, (16)
which generates a formal WKB series for $\omega^{2}$,
$\omega^{2}(\epsilon)=V_{0}-i\epsilon\Bigl(n+\tfrac{1}{2}\Bigr)\sqrt{-2V_{0}^{(2)}}+\sum_{j=2}^{N}c_{j}\,\epsilon^{j}\,$, (17)
with the coefficients $c_{j}=c_{j}\bigl(n,V_{0}^{(2)},V_{0}^{(3)},\ldots,V_{0}^{(2j)}\bigr)$ obtained algebraically from (16).
Because the WKB series is asymptotic, accuracy improves markedly by applying a Padé resummation to the truncated polynomial $P_{N}(\epsilon)\equiv\omega^{2}(\epsilon)$:
$\mathcal{P}_{\tilde{m}/\tilde{n}}(\epsilon)=\frac{\displaystyle\sum_{j=0}^{\tilde{m}}a_{j}\,\epsilon^{j}}{\displaystyle 1+\sum_{k=1}^{\tilde{n}}b_{k}\,\epsilon^{k}}\qquad(\tilde{m}+\tilde{n}=N)\,$, (18)
with $a_{j},b_{k}$ fixed by matching the Taylor expansion of $\mathcal{P}_{\tilde{m}/\tilde{n}}$ to $P_{N}$ up to $\mathcal{O}(\epsilon^{N+1})$. The Padé estimate for the QNM is then
$\omega^{2}=\mathcal{P}_{\tilde{m}/\tilde{n}}(1)\,,\qquad\omega=\sqrt{\mathcal{P}_{\tilde{m}/\tilde{n}}(1)}\,$, (19)
[…]
We use the following steps to find QNMs:
  1. 1.
    Locate the potential maximum $r_{0}$ by solving $V^{\prime}(r_{0})=0$; compute $r_{*0}$ and $V_{0}^{(k)}$ up to the order needed by $N$.
  2. 2.
    Build the WKB quantization condition (15) and extract the truncated series $P_{N}(\epsilon)$ for $\omega^{2}$ as in (17).
  3. 3.
    Construct Padé approximants $\mathcal{P}_{\tilde{m}/\tilde{n}}(\epsilon)$ per (18) and evaluate at $\epsilon=1$ to obtain $\omega$ via (19).
  4. 4.
[…]
The WKB–Padé method is most reliable for moderate to large multipoles and low overtones (heuristically $n\ll\ell$ for single-barrier potentials). Tables I–IV summarize the results obtained with 6th and higher-order WKB methods combined with Padé resummation, while selected modes are validated through time-domain extraction when necessary.
$l_{cr}$ WKB-6 $m=3$ WKB-7 $m=4$ difference
$0.3$ 0.3736420.088936i0.373642-0.088936i 0.3736420.088957i0.373642-0.088957i $0.00536\%$
$0.5$ 0.3736420.088936i0.373642-0.088936i 0.3736590.088971i0.373659-0.088971i $0.0102\%$
$0.7$ 0.3734460.089169i0.373446-0.089169i 0.3737230.088886i0.373723-0.088886i $0.103\%$
$0.9$ 0.3740200.089779i0.374020-0.089779i 0.3732850.088740i0.373285-0.088740i $0.331\%$
$1.13$ 0.3691870.085818i0.369187-0.085818i 0.3691420.085698i0.369142-0.085698i $0.0338\%$
$1.137$ 0.3689990.085650i0.368999-0.085650i 0.3689560.085529i0.368956-0.085529i $0.0341\%$
Table 1: QNMs of the $\ell=2$, $n=0$ axial gravitational perturbations for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
$l_{cr}$ WKB-6 $m=3$ WKB-7 $m=4$ difference
$0.3$ 0.3460990.273533i0.346099-0.273533i 0.3459350.274690i0.345935-0.274690i $0.265\%$
$0.5$ 0.3460990.273533i0.346099-0.273533i 0.3467690.274310i0.346769-0.274310i $0.233\%$
$0.7$ 0.3483650.277361i0.348365-0.277361i 0.3459190.272993i0.345919-0.272993i $1.12\%$
$0.9$ 0.3507470.275420i0.350747-0.275420i 0.3444350.270005i0.344435-0.270005i $1.86\%$
$1.13$ 0.3254870.257860i0.325487-0.257860i 0.3247010.256776i0.324701-0.256776i $0.322\%$
$1.137$ 0.3246270.257215i0.324627-0.257215i 0.3239420.256233i0.323942-0.256233i $0.289\%$
Table 2: QNMs of the $\ell=2$, $n=1$ for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
$l_{cr}$ WKB-6 $m=3$ WKB-7 $m=4$ difference
$0.3$ 0.5994430.092703i0.599443-0.092703i 0.5994430.092703i0.599443-0.092703i $0\%$
$0.5$ 0.5994430.092703i0.599443-0.092703i 0.5994430.092703i0.599443-0.092703i $0\%$
$0.7$ 0.5994650.092675i0.599465-0.092675i 0.5994540.092707i0.599454-0.092707i $0.00547\%$
$0.9$ 0.5988950.092830i0.598895-0.092830i 0.5989870.092770i0.598987-0.092770i $0.0181\%$
$1.13$ 0.5961840.091107i0.596184-0.091107i 0.5962880.091159i0.596288-0.091159i $0.0193\%$
$1.137$ 0.5959270.090979i0.595927-0.090979i 0.5961490.091043i0.596149-0.091043i $0.0384\%$
Table 3: QNMs of the $\ell=3$, $n=0$ for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
$l_{cr}$ WKB-6 $m=3$ WKB-7 $m=4$ difference
$0.3$ 0.5826420.281298i0.582642-0.281298i 0.5826450.281299i0.582645-0.281299i $0.00037\%$
$0.5$ 0.5826420.281298i0.582642-0.281298i 0.5826440.281297i0.582644-0.281297i $0.00026\%$
$0.7$ 0.5826120.280989i0.582612-0.280989i 0.5827750.281350i0.582775-0.281350i $0.0613\%$
$0.9$ 0.5787930.281739i0.578793-0.281739i 0.5797230.281044i0.579723-0.281044i $0.180\%$
$1.13$ 0.5687870.275578i0.568787-0.275578i 0.5679080.274439i0.567908-0.274439i $0.228\%$
$1.137$ 0.5678750.274646i0.567875-0.274646i 0.5673060.274071i0.567306-0.274071i $0.128\%$
Table 4: QNMs of the $\ell=3$, $n=1$ for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
An alternative approach to computing QNMs is based on the direct integration of the perturbation equation in the time domain. The wave-like equation containing the second derivative in time instead of $\omega^{2}$ is discretized on a characteristic grid using the light-cone variables $u=t-r_{*}$ and $v=t+r_{*}$, following the scheme introduced in Gundlach et al. (1994). This method evolves an initial pulse through the effective potential and yields the complete time-dependent signal at a fixed spatial location outside the black hole. At intermediate times, the signal is dominated by exponentially damped oscillations corresponding to the quasinormal ringing.
To extract the dominant frequencies from the numerically obtained waveform, one typically fits the signal to a superposition of damped exponentials. In practice, this is efficiently accomplished with the Prony method, which provides both the real oscillation frequencies and the damping rates. The time-domain approach thus serves as an important cross-check of the WKB–Padé results, and it is particularly useful in regimes where the WKB expansion becomes less reliable. The time-domain integration together with the Prony method has been described in numerous publications (see, for instance, Ishihara et al. (2008); Malik (2025d); Lütfüoğlu (2025c); Churilova et al. (2020); Konoplya and Zhidenko (2014); Qian et al. (2022); Varghese and Kuriakose (2011); Malik (2024b); Konoplya et al. (2020); Dubinsky (2024); Cuyubamba et al. (2016); Bolokhov (2024c); Lütfüoğlu (2025d, e) for details).
For $\ell=2$, the time-domain integration yields
$\omega=0.369359-0.0856186i$
for the least damped mode, while the sixth-order WKB method with Padé approximants gives
$\omega=0.368999-0.085650i.$
The difference between the two methods remains below $0.1\%$, whereas the deviation from the Schwarzschild limit
$\omega_{\text{Schw}}=0.3736715-0.0889625i$
amounts to about $4\%$ for the damping rate and more than $1\%$ for the real oscillation frequency. In other words, for the lowest multipole $\ell=2$, the physical effect is at least one order of magnitude larger than the numerical uncertainty of the WKB approximation (see Figure 3). Moreover, as $\ell$ increases, the WKB error decreases significantly, becoming two or more orders of magnitude smaller than the effect itself (see Figure  4). Therefore, we conclude that the WKB method provides a reliable tool in this context. We can see that once the quantum correction is included, the real oscillation frequency and damping rate are decreased in comparison with their classical (Schwarzschild) limits.
Figure 3: Time-domain profile for $\ell=2$, $l_{cr}=1.137$. The Prony method gives $\omega=0.369359-0.0856186i$ for the fundamental mode $n=0$.
Figure 4: Time-domain profile for $\ell=3$, $l_{cr}=1.137$. The Prony method gives $\omega=0.369373-0.0856553i$ for the fundamental mode $n=0$.
Notice that, by employing the correspondence between QNMs and grey-body factors Konoplya and Zhidenko (2024b); Lütfüoğlu (2025f); Malik (2025e); Bolokhov and Skvortsova (2025d); Han and Gwak (2025), the recently obtained grey-body factors in Dubinsky (2025a) can be reproduced from the numerical data presented in this work.
Taking into account the recent advances in gravitational-wave observations, which have achieved very high signal-to-noise ratios Abac et al. (2025), the estimated $4\%$ effect provides hope that such deviations may become observable. This raises the prospect of distinguishing classical black holes from their regular, quantum-corrected counterparts, even though the geometric modifications are confined primarily to a relatively small region near the event horizon.
V Conclusions
In this work, we have analyzed the gravitational quasinormal spectrum of the Dymnikova black hole, a regular spacetime that can be interpreted either as a phenomenological model with a de Sitter core or as an effective geometry emerging from Asymptotically Safe gravity. Using the WKB method with Padé improvements, supported by time-domain integration, we have obtained the dominant quasinormal frequencies for axial gravitational perturbations. Our results show that the quantum correction parameter $l_{\rm cr}$ systematically modifies the spectrum: both the real oscillation frequencies and the damping rates decrease when the corrections are switched on, so that the ringdown signal becomes longer-lived compared to the Schwarzschild case. In the limit of large $l_{\rm cr}$, the Schwarzschild spectrum is smoothly recovered, confirming the consistency of the model.
It is important to emphasize that the present study has been restricted to the fundamental mode and the first overtone for which the WKB approach is reliable. For higher overtones, however, the situation is expected to be qualitatively different. Since overtones are strongly localized near the event horizon, they should be especially sensitive to the local geometric deformations induced by quantum corrections Konoplya and Zhidenko (2024c); Konoplya et al. (2022). To investigate this regime, one would need to resort to more precise techniques such as the Frobenius (Leaver) method Leaver (1985), which provides convergent series solutions and can capture the full spectrum beyond the eikonal approximation. At the same time, the metric, and consequently the master wave equation, does not have a rational form, so one has to transform or approximate the metric by some rational function for accurate calculations of higher overtones. Having in mind that the fundamental mode and possibly the first overtone are the only realistic candidates for detection in the near future experiments, we leave this analysis, as well as the inclusion of polar perturbations and possible extensions to higher dimensions, for future work.
Acknowledgements.
A. S. is grateful to the Excellence project FoS UHK 2205/2025-2026 and other funds of the FoS UHK for the financial support. B. C. L. and J. R. are grateful to the Excellence project FoS UHK 2205/2025-2026 for the financial support.